# Robust Tail Index Estimation under Random Censoring via Minimum Density Power Divergence Nour Elhouda Guesmia, Abdelhakim Necir∗, Djamel

### v1

@@ Abstract
Based on the minimum density power divergence approach, we propose a robust estimator of the tail index for randomly right-censored data from a Pareto-type distribution. We establish its consistency and asymptotic normality. An extensive simulation study is performed to assess the finite-sample behavior of the estimator in comparison with existing ones. The methodology is further illustrated through an application to a real AIDS survival dataset.
@@ 1. Introduction
Right-censored data pose a common challenge in statistics, as values are only known up to a certain threshold, while those beyond it remain unobserved. This issue frequently arises in fields such as survival analysis, reliability engineering, and insurance. Pareto-type distributions are commonly used to model data exhibiting extreme values. These distributions are characterized by a relatively high probability of producing observations far from the mean, in contrast to the normal distribution. They are crucial in applications like insurance, finance, and environmental studies, where rare events can have a substantial impact. Let be a sample of size from a random variables (rv) and let be another sample from a rv defined on a probability space with continuous cumulative distribution functions (cdf’s) and respectively. Assume that and are independent. Also, we suppose that is right censored by that is each for stage we can only observe the variable and the indicator variable which determines whether or not has been observed. We assume that the tails functions and are regularly varying at infinity (or Pareto-type) of positive tail indices and respectively. On the other terms for any we have
We denote the cdf of by Then, by the independence between and we have which yields that is also regularly varying at infinity, with tail index In the presence of extreme values and right-censored data, several methods have been proposed for estimating the tail index. Numerous techniques have been developed to address the challenges arising from the nature of such data. Many studies have focused on modifying traditional tail index estimation methods, such as Hill’s estimator (Hill, 1975) to make them more suitable for censored data. Einmahl et al. (2008) adapted this estimator to handle the characteristics of right-censored data, and it is given by where denotes the well-known Hill estimator corresponding to the tail index and
stands for an estimator of the proportion of upper non-censored observations
The integer sequence represents the number of top order statistics used in the estimation of the tail index so that and as The sequence of rv’s represents the order statistics pertaining to the sample and denotes the corresponding concomitant values, satisfying for such that On the basis of a Kaplan-Meier integration, Worms and Worms (2014) proposed a consistent estimator to the tail index defined by
where denotes the well-known Kaplan-Meier estimator of cdf (Kaplan and Meier, 1958). The asymptotic normality of is established in Beirlant et al. (2019) by considering Hall’s model (Hall, 1982). The bias reduction in tail index estimation for censored data is addressed by Beirlant et al. (2016), Beirlant et al. (2018) and Beirlant et al. (2019). By using a Nelson-Aalen integration, recently Meraghni et al. (2025) also derived a new estimator for given by
@@
denotes the well-known Nelson-Aalen estimator of cdf (Nelson, 1972). Note that
@@
The remainder of this paper is organized as follows. Section 2 provides a brief overview of the MDPD estimation method, originally introduced by Basu et al. (1998). Section 3 discusses the asymptotic properties of the proposed estimator, namely its consistency and asymptotic normality, with detailed proofs deferred to Section 6. The finite-sample performance of the estimator is examined in Section 5 through a simulation study, including comparisons with existing estimators. Section 7 collects several technical lemmas and propositions. Section 8 presents the figures corresponding to the simulation results.
@@ 2.1. MDPD estimator of for right-censored data
Meraghni et al. (2025) stated that where and are the empirical counterparts of the cdf and the sub-distribution respectively, defined by:
See, for example, Shorack and Wellner (1986) page 294. Here, denotes the left limit of the function at By substituting with the above expression and using the ratio from we obtain:
@@ 3. Consistency and asymptotic normality
Since weak approximations of extreme value theory-based statistics are typically achieved within the second-order framework (see de Haan and Stadtmüller, 1996), it is therefore natural to assume that the cdf satisfies the well-known second-order condition of regular variation. Specifically, we assume that for any

### v3

@@ Abstract
We propose a new robust estimator for the tail index of a Pareto-type distribution under random right-censoring, developed within the framework of the minimum density power divergence. To the best of our knowledge, this is the first application of this divergence to randomly censored extreme value models, providing a novel direction for robust inference in this setting. Under mild regularity conditions, the proposed estimator is shown to be consistent and asymptotically normal. Its finite-sample performance is thoroughly examined through an extensive simulation study, demonstrating enhanced robustness and efficiency compared to existing methods. The practical usefulness of the approach is further illustrated through an application to a real AIDS survival dataset.
@@ 1. Introduction
Right-censored data present a common challenge in statistics, as values are only known up to a certain threshold, while those beyond it remain unobserved. This issue frequently arises in fields such as survival analysis, reliability engineering, and insurance. Pareto-type distributions are commonly used to model data exhibiting extreme values. These distributions are characterized by a relatively high probability of producing observations far from the mean, in contrast to the normal distribution. They are particularly relevant in applications such as insurance, finance, and environmental studies, where rare events can have a substantial impact. Let be a sample of size from a random variables (rv) and let be another sample from a rv both defined on a probability space with continuous cumulative distribution functions (cdf’s) and respectively. Assume that and are independent. Also, we suppose that is right censored by that is each for stage we can only observe the variable and the indicator variable which determines whether or not has been observed. We assume that the tails functions and are regularly varying at infinity (or Pareto-type) of positive tail indices and respectively. On the other terms for any we have
We denote the cdf of by Then, by the independence between and we have which yields that is also regularly varying at infinity, with tail index In the presence of extreme values and right-censored data, several methods have been proposed for estimating the tail index. Numerous techniques have been developed to address the challenges arising from the nature of such data. Many studies have focused on modifying traditional tail index estimation methods, such as Hill’s estimator (Hill, 1975) to make them more suitable for censored data. Einmahl et al. (2008) adapted this estimator to handle the characteristics of right-censored data, and it is given by
where denotes the well-known Hill estimator corresponding to the tail index and stands for an estimator of the proportion of upper non-censored observations The integer sequence represents the number of top order statistics used in the estimation of the tail index so that and as The sequence of rv’s represents the order statistics pertaining to the sample and denotes the corresponding concomitant values, satisfying for such that On the basis of a Kaplan-Meier integration, Worms and Worms (2014) proposed a consistent estimator to the tail index defined by
@@
denotes the well-known Kaplan–Meier estimator of cdf (Kaplan and Meier, 1958). The asymptotic normality of is established in Beirlant et al. (2019) by considering Hall’s model (Hall, 1982). The bias reduction in tail index estimation for censored data is addressed by Beirlant et al. (2016), Beirlant et al. (2018) and Beirlant et al. (2019). By using a Nelson-Aalen integration, recently Meraghni et al. (2025) also derived a new estimator for given by
where
denotes the well–known Nelson–Aalen estimator of cdf (Nelson, 1972). Note that
@@
While robust estimation techniques have been extensively developed for complete data settings, their extension to the context of randomly right-censored observations remains considerably less explored. Such censoring mechanisms are prevalent in survival analysis, reliability theory, and actuarial applications, where incomplete information poses additional statistical challenges. Classical estimators of the extreme value index under censoring—such as Hill-type approaches or those based on Kaplan–Meier and Nelson–Aalen integrals—often exhibit sensitivity to outliers and deviations from the assumed model, thereby limiting their practical robustness. To the best of our knowledge, a comprehensive robust framework for tail index estimation in the presence of censoring and heavy-tailed behavior has not yet been established. In this work, we address this gap by adapting the MDPD methodology to the censored data setting. The resulting estimator preserves the robustness properties inherent to the MDPD approach, while accounting for the incomplete nature of the data and maintaining desirable asymptotic characteristics.
The remainder of this paper is organized as follows. Section 2 provides a brief overview on the MDPD estimation method, originally introduced by Basu et al. (1998). Section 3 discusses the asymptotic properties (consistency and asymptotic normality) of the proposed estimator, with detailed proofs deferred to Section 6. The finite-sample performance of the estimator is examined in Section 4 through a simulation study, including comparisons with existing estimators. Section 5 is devoted to a case study of AIDS survival data. Finally, several technical lemmas and propositions (instrumental to our needs) are compiled in Appendix A, while Appendix B contains the figures related to the simulation results.
@@ 2.1. MDPD estimator of for right-censored data
Meraghni et al. (2025) stated that
where and are the empirical counterparts of the cdf and the sub-distribution respectively, defined by:
See, for example, Shorack and Wellner (1986) page 294. Here, denotes the left limit of the function at Substituting in with its expression above yields
@@ 3. Consistency and asymptotic normality
Since weak approximations of extreme value theory-based statistics are typically achieved within the second-order framework (see, i.e., de Haan and Stadtmüller, 1996), it is therefore natural to assume that the cdf satisfies the well-known second-order condition of regular variation. Specifically, we assume that for any
@@ 5. Real data application
Conclusion
In this work, we addressed the problem of robust estimation of the extreme value index under random right censoring—a setting frequently encountered in survival analysis, reliability engineering, and insurance. While various estimators have been proposed in the literature, including Hill-type methods and those based on Kaplan–Meier or Nelson–Aalen estimators, many of these approaches are sensitive to outliers or rely on strong parametric assumptions.
To overcome these limitations, we introduced a new estimator based on the minimum density power divergence (MDPD), a criterion well known for its robustness in the uncensored context. We carefully extended this methodology to handle randomly right-censored data, accounting for the inherent incompleteness in such settings. To the best of our knowledge, this is the first application of the MDPD framework to tail index estimation under censoring, thereby addressing a notable gap in the robust extreme value literature.
We established the consistency and asymptotic normality of the proposed estimator under standard regular variation assumptions. A comprehensive simulation study demonstrated its strong performance in finite samples, showing an advantageous balance between robustness and efficiency, particularly under model misspecification or contamination.
This contribution opens new perspectives for developing robust methods in censored extreme value models. It also provides a useful foundation for future research, such as adaptive threshold selection, multivariate extensions, and models with dependent censoring.